A fair six-sided die is rolled five times. Find the chance of the event: Exactly one roll is a six.
Answer #1

Probability of rolling a six = 1/6

Let X be the number of outcome as 6 in 5 rolls. Then X ~ Binomial(n = 5, p = 1/6)

Probability of exactly one roll is a six = P(X = 1) = 5C1 * (1/6)1 * (1 - 1/6)5-1

= 5 * (1/6) * (5/6)4

= 0.4018776